OpenStudy (anonymous):

12 1/2 percent is equals in fraction to 3/8 ?? True or False

OpenStudy (anonymous):

1/2=50% 3/8=37,5% False.

OpenStudy (anonymous):

12.5% = 1/8 37.5% =3/8